if f(x)=3x-2 and g(x)=6 -4x find f(x)+ g(x)
Answer by only4christ(12)

Here, you just need to add the two functions.
f(x) = 3x - 2 and g(x) = 6 - 4x, so f(x) + g(x) =
3x - 2 + 6 - 4x......collecting like terms we have
-x + 4..that's it.